find F(2) for this polynomial Fx = -4x^2 + 3x - 1
Answer by nerdybill(7384)

find F(2) for this polynomial Fx = -4x^2 + 3x - 1
.
Simply replace x with a 2:
Fx = -4x^2 + 3x - 1
Fx = -4(2)^2 + 3(2) - 1
Fx = -4(4) + 6 - 1
Fx = -16 + 6 - 1
Fx = -10 - 1
Fx = -11